Given P(x)=2x^2+13x+11 and Q(x)=2x^2+10x+8 Find P(x)+Q(x)
Answer by kamallohia(18)

Just add the coefficients of x^2 , x and the constant terms.
Given P(x)=2x^2+13x+11 and Q(x)=2x^2+10x+8
So, P(x)+Q(x)= (2+2)x^2+(13+10)x+(11+8)
=>P(x)+Q(x)= 4x^2+23x+19.
Hope it's clear.
Kamal Lohia